mhchen:

As n approaches infinity, what is the limit of (2+sin(n^2))/(42+(1/n)) I'll type a cleaner version in the comments

mhchen:

\[\lim_{n \rightarrow \infty}\frac{2+\sin(n^2)}{42+\frac{1}{n}}\]

Vocaloid:

@angle (my calculus is very rusy but iirc the limit of sin(x^2) DNE as x --> infinity b/c the function fluctuates between -1 and 1?)

Angle:

correct

Vocaloid:

so the entire limit would not exist? :S

Angle:

exactly if any part of a limit Does Not Exist (DNE) then the DNE trumps all and the entire limit cannot exist

mhchen:

ah.....doesn't exist then. That's my problem.

mhchen:

Wait so if I'm trying to compare 2 functions on which one has a 'higher dominance', like little-o. Say function f(x) and g(x). I'd do \[\lim_{x \rightarrow \infty}\frac{f(x)}{g(x)}\] If it becomes 0, g(x) dominates. If it becomes infinity, f(x) dominates, if it's a constant, they have the same theta-notation. If they don't exist....would they also have the same theta notation?

Angle:

I think that it might be that the function that causes the DNE would dominate

mhchen:

hmm, that would make more sense